40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 276 of 335  Not logged in ELOG logo
ID Date Author Type Category Subject
  2997   Thu May 27 02:22:24 2010 kiwamuUpdateGreen Lockingmore details

 Here are some more plots and pictures about the end PDH locking with the green beam. 

-- DC reflection

 I expected that the fluctuation of the DC reflection had 1% from the resonant state to the anti-resonant state due to its very low finesse.

This values are calculated from the reflectivity of ETM measured by Mott before (see the wiki).

In my measurement I obtained  DC reflection of V_max=1.42 , V_min=1.30  at just after the PD.

These numbers correspond to 7.1% fluctuation. It's bigger than the expectation.

I am not sure about the reason, but it might happen by the angular motion of test masses (?)

 

--- time series

Here is a time series plot. It starts from openloop state (i.e. feedback disconnected).

At t=0 sec I connected a cable which goes to the laser pzt, so now the loop is closed.

You can see the DC reflection slightly decreased and stayed lower after the connection.

The bottom plot represents the feedback signal measured before a sum amp. which directly drives the pzt.

stimes.png

 

 

-- length fluctuation  

One of the important quantities in the green locking scheme is the length fluctuation of the cavity.

It gives us how much the frequency of the green beam can be stabilized by the cavity. And finally it will determine the difficulty of PLL with the PSL.

I measured a spectrum of the pzt driving voltage [V/Hz1/2] and then converted it to a frequency spectrum [Hz/Hz1/2].

I used the actuation efficiency of 1MHz/V for the calibration, this number is based on the past measurement.

spectrum.png

RMS which is integrated down to 1Hz  is 1.6MHz.

This number is almost what I expected assuming the cavity swings with displacement of x ~< 1um.

 

-- flashing

A picture below is a ETMx CCD monitor.

One of the spot red circled in the picture blinks when it's unlocked. And once we get the lock the spot stays bright.

ETMX_small.png

 

  2996   Wed May 26 22:22:03 2010 AlbertoConfiguration40m UpgradingArm cavity length

The second sideband is resonant in the arms for a cavity length of 37.9299m.

The nearest antiresonant arm lengths for f2 (55MHz) are 36.5753m and 39.2845m.

If we don't touch the ITMs, and we use the room we still have now on the end tables, we can get to 37.5m.

This is how the power spectrum at REFL would look like for perfect antiresonance:

reflRFpowerVsArmLength_3658m.png

And this is how it looks like for 37.5m:

reflRFpowerVsArmLength_3750m.png

Or, god forbid, we change the modulation frequencies...

  2995   Wed May 26 18:54:55 2010 AidanSummaryGreen LockingMounted Crystal 724 in the Doubling Oven

Andri and I mounted the Raicol Crystal #724 in one of the new Covesion Ovens. The procedure was the same as before - see elog entry here.

There was one issue - the glass plate that goes on top of the crystal is coated on one side with ITO (Indium-Tin Oxide) and it's not 100% certain that this was mounted in the correct orientation. It is virtually impossible to tell which side of the glass is coated.

The base plate of the oven was tapped for an M3 hole. We retapped it for an 8-32 and bolted it to a post and that one of the New Focus 4-axis translation stage. The assembly is currently bolted to the PSL table, awaiting use.

  2994   Wed May 26 17:10:09 2010 AlbertoUpdate40m UpgradingRF Generation box

This is how the RF generation box might soon look like:

Visio-frequencyGenerationBox_wiringSchematic.png

A dedicated wiki page shows the state of the work:

http://lhocds.ligo-wa.caltech.edu:8000/40m/Upgrade_09/RF_System/frequency_generation_box#preview

  2993   Wed May 26 15:34:00 2010 JenneUpdateIOOMirrors moved in prep for round 2 of MC mode measuring

Quote:

That's true. But I thought that you measured the mode after those optics and the effect of them is already included.

So:

  • We need to model the transmissive optics in order to understand the measured mode which is different from the MC mode slightly.
  • We just can calculate the modes based on the measurement in order to figure out the realistic positions of the MMT1 and MMT2.

Quote:

Rana pointed out that the anticipated mode calculation should be modified to include the index of refraction of the crystals in the Faraday, and the polarizers in the Faraday.  This may affect where we should put MMT1, and so this should be completed before round 2 measurements are taken, so that we can move MMT1.

 

 Yes, the measured mode takes all of this into account.  But in Kevin's plot, where he compares 'measured' to 'expected', the expected doesn't take the Faraday optics into account.  So I should recalculate things to check how far off our measurement was from what we should expect, if I take the Faraday into account.  But for moving forward with things, I can just use the mode that we measured, to adjust (if necessary) the positions of MMT1 and MMT2.  All of the other transmissive optics (that I'm aware of) have already been included, such as the PRM and the BS.  This included already the air-glass curved interface on the PRM, etc.

  2992   Wed May 26 14:38:02 2010 KojiUpdateGreen Lockinglocked

Congratulation! Probably you are right, but I could not get this is a real lock or something else.

1) How much was the fringe amplitude (DC) of the reflected beam? (Vref_max=XXX [V] and Vref_min=YYY [V])
    Does this agree with the expectation?

2) Do you have the time series? (V_ref and V_error)

Quote:

I guess I succeeded in locking of the cavity with the green beam 

 Strictly speaking, the laser frequency of the end NPRO is locked to the 40 meter arm cavity.

Pictures, some more quantitative numbers and some plots are going to be posted later.

 


After the alignment of the cavity I could see DC fringes in its reflection. Also I could see the cavity flashing on the monitor of  ETMY_CCD.

I drove the pzt of the NPRO with f=200kHz, and then the spectrum analyzer showed 200kHz beat note in the reflection signal. This means it's ready to PDH technique.

And then I made a servo loop with two SR560s, one for a filter and the other for a sum amp.

After playing with the value of the gain and the sign of the feedback signal, the laser successfully got lock. 

 

To make sure it is really locked, I measured the open loop transfer function of the PDH servo while it stayed locked. The result is shown in the attached figure.

The measured data almost agrees with the expected curve below 1kHz, so I conclude it is really locked.

However the plot looks very noisy because I could not inject a big excitation signal into the loop. If I put a big excitation, the servo was unlocked.

The current servo is obviously too naive and it only has f-1 shape, so the filter should be replaced by a dedicated PDH box as we planed.

 

  2991   Wed May 26 14:28:01 2010 KojiUpdateIOOMirrors moved in prep for round 2 of MC mode measuring

That's true. But I thought that you measured the mode after those optics and the effect of them is already included.

So:

  • We need to model the transmissive optics in order to understand the measured mode which is different from the MC mode slightly.
  • We just can calculate the modes based on the measurement in order to figure out the realistic positions of the MMT1 and MMT2.

Quote:

Rana pointed out that the anticipated mode calculation should be modified to include the index of refraction of the crystals in the Faraday, and the polarizers in the Faraday.  This may affect where we should put MMT1, and so this should be completed before round 2 measurements are taken, so that we can move MMT1.

 

  2990   Wed May 26 12:59:26 2010 josephbUpdateCDSCreated sus, sup, scx, spx models

I created the sus model, which is the suspension controller for ITMX, ITMY, BS, PRM, SRM.  I also created sup, which is the suspension plant model for those same optics.

Updated /cvs/cds/caltech/target/fb  master and daqdrc files to add SUS, SUP models.  Megatron's /etc/rc.d/rc.local file has been updated to include all the necessary models as well.

The suspension controller needs the Binary IO outputs need to be checked and corrected if wrong by changing the constant connected to the exclusive or gates. Right now its using the end suspension binary output values which may not be correct.

  2989   Wed May 26 10:58:29 2010 josephbUpdateCDSNew RCG checkout for use with all machines plus some issues

Now that we have multiple machines we'd like to run the new front end code on, I'm finding it annoying to have to constantly copy files back and forth to have the latest models on different machines.  So I've come to the conclusion that Rana was right all along, and I should working somewhere in /cvs/cds/caltech which gets mounted by everyone. 

However, this leads to the svn problem: I.e. I need recent code checked out from the RCG repository, but our current /cvs/cds/caltech/cds/advLigo directory is covered by the 40m SVN.  So for the moment, I've checked out the advLigoRTS from https://redoubt.ligo-wa.caltech.edu/svn/advLigoRTS/trunk into /cvs/cds/caltech/cds/advLigoRTS.  This directory will be kept as up to date as I can keep it, both by running svn update to get Alex/Rolf's changes and on my end by keeping the new and updated models.  It will remain linked the RCG repository and not the 40m repository.  At some point a better solution is needed, but its the best I can come up with for now.

Also, because we are starting to compile on different machines sometimes, you may run into a problem where a code won't run on a different machine.  This can be fixed by commenting out some lines in the startup script.  Go to the /cvs/cds/caltech/scripts directory.  Then edit the associated startSYS file by commenting out the lines that look like:

if [ `hostname` != megatron ]; then
echo Cannot run `basename $0` on `hostname` computer
exit 1
fi

Unfortunately, this gets reverted each time "make SYS" and "make install-SYS" gets run.

The other issue this leads to is that some machines don't have as many CPUs available as others.  For example our new thin 1U machines have only 4 dual cores (8 CPUs total).  This means the specific_cpu setting of any of the codes cannot be higher than 7 (cores being numbered 0 through 7).  Core 0 is reserved for the real time kernel, and Core 1 will be used on all machines for the IO processor. This leaves only cores 2 through 7 available for models to use which include LSC, LSP, SUS, SUP, SPY, SCY, SPX, SCX, OMC, OMP, OAF, OAP?, IOC, IOP.  Since there are more than 6 models, duplication in final production code of specific_cpus will be necessary.  Codes which are all running on Megatron at one point will have to be rebuilt with new specific_cpu values when run on the actual final machine.

  2988   Wed May 26 04:14:21 2010 kiwamuUpdateGreen Lockinglocked

I guess I succeeded in locking of the cavity with the green beam 

 Strictly speaking, the laser frequency of the end NPRO is locked to the 40 meter arm cavity.

Pictures, some more quantitative numbers and some plots are going to be posted later.

 


After the alignment of the cavity I could see DC fringes in its reflection. Also I could see the cavity flashing on the monitor of  ETMY_CCD.

I drove the pzt of the NPRO with f=200kHz, and then the spectrum analyzer showed 200kHz beat note in the reflection signal. This means it's ready to PDH technique.

And then I made a servo loop with two SR560s, one for a filter and the other for a sum amp.

After playing with the value of the gain and the sign of the feedback signal, the laser successfully got lock. 

 

To make sure it is really locked, I measured the open loop transfer function of the PDH servo while it stayed locked. The result is shown in the attached figure.

The measured data almost agrees with the expected curve below 1kHz, so I conclude it is really locked.

However the plot looks very noisy because I could not inject a big excitation signal into the loop. If I put a big excitation, the servo was unlocked.

The current servo is obviously too naive and it only has f-1 shape, so the filter should be replaced by a dedicated PDH box as we planed.

Attachment 1: OLTF_endPDH.png
OLTF_endPDH.png
  2987   Wed May 26 00:50:16 2010 JenneUpdateIOOMirrors moved in prep for round 2 of MC mode measuring

[Jenne, Kevin, Kiwamu]

We moved some optics in preparation for measuring the MC mode after the first MMT curved optic, RoC -5m. 

Kevin and I found the box of DLC (sp?) mounts with the 2" Y1-45P optics in the clean tupperware boxes.  We removed one of the Y1-45P's, and replaced it with the MMT1 -5m optic, which was baked several weeks ago.  We left the Y1-45P on the cleanroom table next to where the MMT optics are.  We placed this MMT mirror in the place it belongs, according to Koji's table layout of the BS table. 

We drag wiped one of the other Y1-45P's that was in the box since it was dirty, and then placed the optic on the IOO table, on the edge closest to the BS table, with the HR side facing the BS table, so that the beam reflected off the curved mirror is reflected back in the direction of the BS table.  This was aligned so the beam hits the same PZT mirror we were using last time, to get the beam out of the BS chamber door.  We left a razor dump on the edge of the BS table, by the door, which will need to be removed before actual measurements can take place. 

Rana pointed out that the anticipated mode calculation should be modified to include the index of refraction of the crystals in the Faraday, and the polarizers in the Faraday.  This may affect where we should put MMT1, and so this should be completed before round 2 measurements are taken, so that we can move MMT1.

Also, the optics are in place now, and the beam is going out the BS chamber door, but we have not yet measured distances (design distances quoted on the MMT wiki page), and confirmed that everything is in the right place.  So there is a bit more work required before beginning to measure round 2.

 

Note:  While I was poking around on the BS table, I had to move several optics so that we could fit MMT1 in the correct place.  When preparing to move these optics, I found 2 or 3 that were totally unclamped. This seems really bad, especially for tall skinny things which can fall over if we have an earthquake.  Even if something is in place temporarily, please clamp it down.

  2986   Tue May 25 17:22:56 2010 KevinUpdateIOOBeam Profile After Mode Cleaner

Quote:

Very nice as usual. Can you add the curve to show the ideal mode of the MC on the profile plot?

Quote:

I fit the data from the beam profile that Jenne measured on 5/21/2010. The distances are measured from halfway between MC1 and MC3 to the beam scanner. The fits give the following where w0 is the waist size and z0 is the distance from the waist to halfway between MC1 and MC3.

For the horizontal profile:

reduced chi^2 = 0.88

z0 = (1 ± 29) mm

w0 = (1.51 ± 0.01) mm

For the vertical profile:

reduced chi^2 = 0.94

z0 = (673 ± 28) mm

w0 = (1.59 ± 0.01) mm

I calculated the radius of curvature of MC2 using these values of w0:

horizontal: (16.89 ± 0.06) m

vertical:   (17.66 ± 0.07) m

For this calculation, I used the value of (13.546 ± .0005) m for the length of the mode cleaner measured on 6/10/2009. The specification for the radius of curvature of MC2 is (18.4 ± 0.1) m.

Here is the plot with the ideal mode of the mode cleaner shown in brown. The ideal mode was plotted with the radius of curvature of 18.4. The blue curve is the fit to the vertical beam radius, the purple curve is the fit to the horizontal beam radius, * denotes a data point from the vertical data, and + denotes a data point from the horizontal data.

Attachment 1: mcfit.png
mcfit.png
  2985   Tue May 25 17:09:22 2010 KojiUpdateIOOBeam Profile After Mode Cleaner

Very nice as usual. Can you add the curve to show the ideal mode of the MC on the profile plot?

Quote:

I fit the data from the beam profile that Jenne measured on 5/21/2010. The distances are measured from halfway between MC1 and MC3 to the beam scanner. The fits give the following where w0 is the waist size and z0 is the distance from the waist to halfway between MC1 and MC3.

For the horizontal profile:

reduced chi^2 = 0.88

z0 = (1 ± 29) mm

w0 = (1.51 ± 0.01) mm

For the vertical profile:

reduced chi^2 = 0.94

z0 = (673 ± 28) mm

w0 = (1.59 ± 0.01) mm

I calculated the radius of curvature of MC2 using these values of w0:

horizontal: (16.89 ± 0.06) m

vertical:   (17.66 ± 0.07) m

For this calculation, I used the value of (13.546 ± .0005) m for the length of the mode cleaner measured on 6/10/2009. The specification for the radius of curvature of MC2 is (18.4 ± 0.1) m.

 

  2984   Tue May 25 17:04:37 2010 KevinUpdate Beam Profile After Mode Cleaner

I fit the data from the beam profile that Jenne measured on 5/21/2010. The distances are measured from halfway between MC1 and MC3 to the beam scanner. The fits give the following where w0 is the waist size and z0 is the distance from the waist to halfway between MC1 and MC3.

For the horizontal profile:

reduced chi^2 = 0.88

z0 = (1 ± 29) mm

w0 = (1.51 ± 0.01) mm

For the vertical profile:

reduced chi^2 = 0.94

z0 = (673 ± 28) mm

w0 = (1.59 ± 0.01) mm

I calculated the radius of curvature of MC2 using these values of w0:

horizontal: (16.89 ± 0.06) m

vertical:   (17.66 ± 0.07) m

For this calculation, I used the value of (13.546 ± .0005) m for the length of the mode cleaner measured on 6/10/2009. The specification for the radius of curvature of MC2 is (18.4 ± 0.1) m.

In the following plots, the blue curve is the fit to the vertical beam radius, the purple curve is the fit to the horizontal beam radius, * denotes a data point from the vertical data, and + denotes a data point from the horizontal data.

Attachment 1: mcfit.png
mcfit.png
Attachment 2: mcerrors.png
mcerrors.png
  2983   Tue May 25 16:40:27 2010 josephb, alexUpdateCDSFinally tracked down why new models wouldn't talk to each other

The problem with the new models using the new shared memory/dolphin/RFM defined as names in a single .ipc file.

The first is the no_oversampling flag should not be used.  Since we have a single IO processor handling ADCs and DACs at 64k, while the models run at 16k, there is some oversampling occuring.  This was causing problems syncing between the models and the IOP.

It also didn't help I had a typo in two channels which I happened to use as a test case to confirm they were talking.  However, that has been fixed.

  2982   Tue May 25 16:32:26 2010 kiwamuHowToElectronicsfront ends are back

 [Alex, Joe, Kiwamu]

Eventually all the front end computers came back !! 

There were two problems.

(1): C0DCU1 didn't want to come back to the network. After we did several things it turned the ADC board for C0DCU1 didn't work correctly.

(2): C1PEM1 and C0DAQAWG were cross-talking via the back panel of the crate.


(what we did)

* installed a VME crate with single back panel to 1Y6 and mounted C1PEM1 and C0DAQAWG on it. However it turned out this configuration was bad because the two CPUs could cross-talk via the back panel.

* removed the VME crate and then installed another VME crate which has two back panels so that we can electrically separate C1PEM1 and C0DAQAWG.  After this work, C0DAQAWG started working successfully.

 * rebooted all the front ends, fb40m and c1dcuepics.

 * reset the RFM bypath. But these things didn't bring C0DCU1 back.

 * telnet to C0DCU1 and ran "./startup.cmd" manually. In fact "./startup.cmd" should automatically be called when it boots.

 * saw the error messages from "./startup.cmd" and found it failed when initialization of the ADC board. It saids "Init Failure !! could not find ICS"

*  went to 1Y7 rack and checked the ADC. We found C0DCU1 had two ADC boards, one of two was not in used.

* disconnected all two ADCs and put back one which had not been in used. At the same time we changed the switching address of this ADC to have the same address as the other ADC. 

* powered off/on 1Y7 rack. Finally C0DCU1 got back.

* burtrestored the epics to the last Friday, May 21st 6:07am

  2981   Tue May 25 10:06:09 2010 kiwamuUpdateElectronicsbad power supply of a vme rack

 I got a VME crate from Peter's lab. It is already installed in 1Y6 instead of the old broken one.

I checked its power supply, and it looked fine. It successfully supplies +5, +12 and -12 V. And then I put c0daqawg and c1pem1 back from 1Y7.

Now I am trying to reboot all the front end computers with Peter's VME crate. A picture of the VME crate will be updated later.

  2980   Tue May 25 09:12:46 2010 kiwamuConfigurationGreen Lockingeffect from air conditioner

We should completely turn off the air conditioner when working on green locking.

Even if green beams propagates inside of chambers, the air conditioner does affect the spatial jitter of the beam. 

The attached picture was taken when Steve and I were seeing how the green beam jittered. 

At that time the beam was injected from the end table and going through inside of the ETM, the ITM and the BS camber.

Eventually it came out from the camber and hit the wall outside of the chamber. It was obvious, we could see the jittering when the air cond. was ON.

Attachment 1: green_spot.png
green_spot.png
  2979   Tue May 25 07:58:23 2010 kiwamuUpdateelogelog down

I found the elog got down around 7:30 am in this morning.

So I restarted it by running the script: "start-elog-nodus" as instructed on the wiki.

http://lhocds.ligo-wa.caltech.edu:8000/40m/How_To/Restart_the_elog

  2978   Tue May 25 07:22:59 2010 kiwamuUpdateElectronicsbad power supply of a vme rack

Notes on May 25th

 Don't do the following things !! This causes bad cross-talking of CPUs mounted on the crate.

 


I moved c0daqawg and c1pem1 from 1Y6 vme crate to 1Y7 crate due to the bad power supply.

Another problem: c0dcu1 doesn't come back to the network. 

After moving them, I tried to get back them into the RFM network. However  c0dcu1 never came back, it still indicates red in C0DAQ_DETAIL.adl screen.

Alberto and I did even "nuclear option" (as instructed), but no luck.

  2977   Tue May 25 00:10:24 2010 ZachUpdateIOOIMC table leveled again

The IMC table had to be leveled again, for two reasons: 1) It was un-leveled when Jenne and Kiwamu removed an extra beam dump when they took the beam profile measurements, and 2) the stack of weights I had put there before was too tall to allow the beam to pass (I didn't realize that the BS chamber is offset a bit to the north, so the beam passes right over the NE edge of the IMC table).

First of all, I was wrong before when I said that the stack of weights was 4 blocks tall; it was 6 blocks tall. I re-leveled the table this afternoon by removing the top three blocks and placing them immediately south of the bottom three in the original stack, while also moving the circular weight north of its previous position. The table is now balanced roughly to within the tolerance of the bubble level I was using.

After the leveling, I tried to re-lock the modecleaner. Upon removing the beam block on the PSL table, I got some sort of resonance flashes on the MC TRANS monitor. With some minor adjustments to MC2&3, I was able to get a decent TEM00 mode to hit. The cavity wouldn't lock, so I went to the AP table and checked to make sure that the REFL beam was hitting the PD. It was, but the beam was very close to the edge of the focusing lens, so I moved the steering mirror slightly to make the situation a little better.

I then went to the control room to finish the by-now-mundane task of fine-tuning the MC lock, but today's was a worthier opponent. For some reason, the thing didn't want to lock for more than a few seconds at a time. I saw that the spot on MC2 was quite a bit off-center, so I ran the MC2_spot_xxx scripts to get it visually in place, then revisited the AP table to ensure that the REFL beam was still on the PD. No dice.

I don't know what was different. I had Ameristat over the opening between the tanks, with posters on top and on the sides (as usual), and I checked to ensure that the servo gains were at the appropriate levels. Joe pointed out that IOO VME was not responding, but we didn't seem to think that this was the problem (based on nothing I can put in words or stick figure cartoons), and the "alive" indicator on the Auto-Lock control in the MEDM screen was not blinking, as it usually is, but I don't know what bearing this has on anything.

I will try to lock again tomorrow.

 

  2976   Mon May 24 16:34:22 2010 KevinUpdatePSLND Filters for 2W Beam Profile

I tried to measure the beam profile of the 2W laser today but ran into problems with the ND filters. With the measurements I made a few weeks ago, I used a reflective ND 4.0 filter on the PD. The PD started to saturate and Koji and I noticed that a lot of the metallic coating on the filter had been burnt off. Koji told me to use an absorptive ND 4.0 filter in front of a reflective ND 0.6 filter. I tried this today but noticed that a few holes were being burned into the absorptive filter and that the coating on the reflective filter behind it was also being burned off in spots. I don't think we wanted to use a polarizing beam splitter to reduce the power before the PD but I didn't want to ruin any more filters.

  2975   Mon May 24 14:28:35 2010 kiwamuUpdateElectronicsbad power supply of a vme rack

In this morning I found daqawg didn't work.

After looking for the cause, I found one of the vme racks mounted on 1Y6 doesn't work correctly.

It looks like the vme rack mounting c0daqawg could not supply any power to the frontends.

 

Now Steve and I are trying to look for a spare for it.

  2974   Mon May 24 11:32:05 2010 ranaUpdateLSC40mUpgrade Field Power and RF Power Spectrum at the ports. 38m/38.55m arm length issue.

Quote:

 

 If you have a working 40m Optickle model, put it in a common place in the SVN, not in your own folder.

I can't figure out why changing the arm length would effect the RF sidebands levels. If you are getting RF sidebands resonating in the arms, then some parameter is not set correctly.

As the RF sideband frequency gets closer to resonating in the arm, the CARM/DARM cross-coupling to the short DOFs probably gets bigger.

I uploaded the latest iscmodeling package to the SVN under /trunk. It includes my addition of the 40m Upgrade model: /trunk/iscmodeling/looptickle/config40m/opt40mUpgrade2010.m.

I don't know the causes of this supposed resonances yet. I'm working  to try to understand that. It would be interesting also to evaluate the results of absolute length measurements.

Here is what I also found:

reflRFpowerVsArmLength.png

It seems that 44, 66 and 110 are resonating.

If that is real, than 37.5m could be a better place. Although I don't have a definition of "better" yet.  All I can say is these resonances are smaller there.

  2973   Mon May 24 10:03:14 2010 ranaUpdateLSC40mUpgrade Field Power and RF Power Spectrum at the ports. 38m/38.55m arm length issue.

 

 If you have a working 40m Optickle model, put it in a common place in the SVN, not in your own folder.

I can't figure out why changing the arm length would effect the RF sidebands levels. If you are getting RF sidebands resonating in the arms, then some parameter is not set correctly.

As the RF sideband frequency gets closer to resonating in the arm, the CARM/DARM cross-coupling to the short DOFs probably gets bigger.

  2972   Mon May 24 07:53:57 2010 steveConfigurationPEMair cond. just turned ON

IFO room temp 27.5C , Please remember to turn AC back on !

Attachment 1: acoff.jpg
acoff.jpg
  2971   Fri May 21 16:41:38 2010 Alberto, JoUpdateComputersIt's a boy!

Today the new Dell computer for the GSCS (General SURF Computing Side) arrived.

We put it together and hooked it up to a monitor. And guess what? It works!

I'm totally impressed by how the Windows get blurred on Windows 7 when you move them around. Good job Microsoft! Totally worth 5 years of R&D.

  2970   Fri May 21 16:37:35 2010 steveConfigurationVACtp3's drypump replaced

Bob replaced the tipseal in an other drypump and I swapped it in. TP3 turbo is running again, it's foreline pressure is 40 mTorr. The RGA is still off

  2969   Fri May 21 16:27:45 2010 AlbertoOmnistructureEnvironmentThe control room is molding...

... not just because we haven't locked the interferometer for quite some time. I mean, it literally stinks. The chiller's chiller is molding. Its' dripping water and there's mold all under it (Jo just confirmed: "yeah, it's mold").

Someone from Caltech maintenance just crossed the door. Hopefully he'll help us fix it.

I'll keep you updated. Stay tuned.

  2968   Fri May 21 16:24:11 2010 KojiUpdateLSC40mUpgrade Field Power and RF Power Spectrum at the ports. 38m/38.55m arm length issue.

1. Give us the designed arm length. What is the criteria?

2. The arm lengths got shorter as the ITMs had to shift to the end. To make them longer is difficult. Try possible shorter length.

  2967   Fri May 21 14:31:46 2010 josephb,alexUpdateCDSNew computer and IO chassis working in 1X4

Alex brought over a ADC, DAC, and PCIe card which goes into the computer and talk to the IO chassis.  We tried installing the new "small" IO chassis in 1X4, but initially it couldn't find the ADC/DAC boards, just the Contec Binary out board.

We tried several different configurations (different computer, different IO chassis, the megatron chassis, the megatron IO chassis with new cards, a new IO chassis with megatron cards.

The two things were concluded once we got a new IO chassis talking with the new computer. 

1) Its possible one of the slots is in the IO chassis as we didn't see the ADC until we moved it to a different slot in the new chassis

2) The card that Alex had brought over to put in the computer doesn't behave well with these IO chassis.  He went back to downs to try and figure out what it is, and test it with the chassis over there.

Currently, Megatron's IO chassis is sitting on a table behind racks 1Y5 and 1Y6, along with the new "large" IO chassis.  Megatron's PCIe card for talking to the IO chassis was moved to the computer in 1X4.  The computer in 1X4 is currently being called c1iscex with IP 192.168.113.80. 

  2966   Fri May 21 11:56:34 2010 AlbertoUpdate40m Upgrading40mUpgrade Field Power and RF Power Spectrum at the ports. 38m/38.55m arm length issue.

I update my old 40mUpgrade Optickle model, by adding the latest updates in the optical layout (mirror distances, main optics transmissivities, folding mirror transmissivities, etc). I also cleaned it from a lot of useless, Advanced LIGO features.

I calculated the expected power in the fields present at the main ports of the interferometer.

I repeated the calculations for both the arms-locked/arms-unlocked configurations. I used a new set of functions that I wrote which let me evaluate the field power and RF power anywhere in the IFO. (all in my SVN directory)

As in Koji's optical layout, I set the arm length to 38m and I found that at the SP port there was much more power that I woud expect at 44Mhz and 110 MHz.

It's not straightforward to identify unequivocally what is causing it (I have about 100 frequencies going around in the IFO), but presumably the measured power at 44MHz was from the beat between f1 an f2 (55-11=44MHz), and that at 110MHz was from the f2 first sidebands.

Here's what i found:

RFPower_locked_38m.png

RFPower_unlocked_38m.png

FieldPower_locked_38m.png

FieldPower_unlocked_38m.png

 

I found that When I set the arm length to 38.55m (the old 40m average arm length), the power at 44 and 110 MHz went significantly down. See here:

RFPower_3855m.png

 FieldPower_3855m.png

I checked the distances between all the frequencies circulating in the IFO from the closest arm resonance to them.

I found that the f2 and 2*f2 are two of the closest frequencies to the arm resonance (~80KHz). With a arm cavity finesse of 450, that shouldn't be a problem, though.

40mUpgrade_distanceFromResonance_38m.png

 I'll keep using the numbers I got to nail down the culprit.

Anyways, now the question is: what is the design length of the arms? Because if it is really 38m rather than 38.55m, then maybe we should change it back to the old values.

  2965   Fri May 21 08:28:29 2010 steveConfigurationVACTP3 turned off

Quote:

[Jenne, Kiwamu, and Steve via phone]

Around 9:30pm, Kiwamu and I came back from dinner, and were getting ready to begin the beam scan measurements.  I noticed that one of the vacuum pumps was being very loud.  Kiwamu noted that it is the fore pump for TP3's turbo, which he and Steve replaced in January (elog 2538).  We had not noticed these noises before leaving for dinner, around 8pm.

We called Steve at home, and he could hear the noise through the phone. He said that even though it was really loud, since it was reading 3.3mTorr (on the display of the controller, in the vacuum rack just above head-height) which is close to the nominal value, it should be fine to leave.  He will check it out in the morning.  If it had been reading at or above ~1Torr, that's indicative of it being really bad, and we would have needed to shut it off.

For future reference, in case we need to turn it off, Steve said to use the following procedure:

1. Close VM3, to isolate the RGA, which is what this pump is currently (while we're at atmosphere) pumping on.  I don't know if there are other things which would need to be shut at this stage, if we were at vacuum nominal.

2. Close VM5, which is right in front of TP3, so TP3's pump is just pumping on itself.

3. Push the "Stop" button on the Turbo controller for TP3, in the vacuum rack, about waist level.  Turning off the turbo will also turn off the fore pump.

UPDATE, 1am: The controller in the rack is reading 3.1mTorr, so the pump, while still noisy, still seems to be working.

 The foreline pressure of TP3 is 2.9 mTorr  The drypump is loosing it's bearing and it is very noisy.

V3, V5 closed and TP3 small turbo controller off. This turned off the noisy forepump that has to be replaced.

RGA is running at cc4 2e-6 Torr

The RGA was turned off at cc4 1e-5 Torr

  2964   Fri May 21 00:51:06 2010 JenneUpdateIOOFirst MC mode measuring (hopefully) done

[Jenne, Kiwamu, Steve]

Round 1 of measuring the MC mode is pretty much done.  Yay.

Earlier today, Steve and I launched the MC beam off the flat mirror just after the Faraday, and sent it down toward ETMY(new convention). We ended up not being able to see it all the way at the ETM because we were hitting the beam tube, but at the ITM chamber we could see that the beam looked nice and circle-y, so wasn't being clipped in the Faraday or anywhere else.  To do this we removed 2 1inch oplev optics.  One was removed from the BS table, and wrapped in foil and put in a plastic box.  The other was just layed on its' side on the BS table. 

I then took the beam out of the BS chamber, in order to begin measuring the mode.  I left the flat fixed mirror in the place of what will be PZT SM1, and instead used the PZT mirror to turn the beam and get it out the BS chamber door.  (Thoughts of getting the beam to the BS oplev table were abandoned since this was way easier, since Kiwamu and Steve had made the nifty table leg things.)  Kiwamu and I borrowed an 2inch 45P Y1 optic from the collection on Koji's desk (since we have ZERO 2inch optics on the random-optic-shelf....no good), to shoot the beam down the hallway of the Yarm (new convention).  We used the beam scan on a rolling cart to measure the beam at various distances.  I made some sweet impromptu plum bobs to help make our distance measurements a bit more accurate.

We stopped at ~25 feet from the BS chamber, since the spot was getting too big for the beam scanner.  If it turns out that I can't get a good fit with the points I have, I'll keep everything in-chamber the same, and do the farther distances using the good ol' razor blade technique.

I have measurements for the distances between the beam scan head and the opening of the BS chamber.  Tomorrow, or very soon after, I need to measure the distances in-chamber between the MC and the BS chamber opening.  Plots etc will come after I have those distances.

Next on the to-do list:

1.  Measure distances in-chamber for first mode scan.

2.  Plot spot size vs. distance, see if we need more points. Take more points if needed.

3.  Put in MMT1, repeat measurements.

4.  Put in MMT2, rinse and repeat.

5.  Move the PZT mirror to its new place as SM1, and figure out how to connect it.  Right now the little wires are hooked up on the BS table, but we're going to need to make / find a connector to the outside world from the IOO table. This is potentially a pretty big pain, if we don't by happenstance have open connectors on the IOO table.

  2963   Fri May 21 00:30:44 2010 JenneUpdateVACTP3 fore pump is very loud

[Jenne, Kiwamu, and Steve via phone]

Around 9:30pm, Kiwamu and I came back from dinner, and were getting ready to begin the beam scan measurements.  I noticed that one of the vacuum pumps was being very loud.  Kiwamu noted that it is the fore pump for TP3's turbo, which he and Steve replaced in January (elog 2538).  We had not noticed these noises before leaving for dinner, around 8pm.

We called Steve at home, and he could hear the noise through the phone. He said that even though it was really loud, since it was reading 3.3mTorr (on the display of the controller, in the vacuum rack just above head-height) which is close to the nominal value, it should be fine to leave.  He will check it out in the morning.  If it had been reading at or above ~1Torr, that's indicative of it being really bad, and we would have needed to shut it off.

For future reference, in case we need to turn it off, Steve said to use the following procedure:

1. Close VM3, to isolate the RGA, which is what this pump is currently (while we're at atmosphere) pumping on.  I don't know if there are other things which would need to be shut at this stage, if we were at vacuum nominal.

2. Close VM5, which is right in front of TP3, so TP3's pump is just pumping on itself.

3. Push the "Stop" button on the Turbo controller for TP3, in the vacuum rack, about waist level.  Turning off the turbo will also turn off the fore pump.

UPDATE, 1am: The controller in the rack is reading 3.1mTorr, so the pump, while still noisy, still seems to be working.

  2962   Fri May 21 00:21:24 2010 JenneUpdateSAFETYDon't walk down the Y(new convention) arm tonight!

All clear.

Quote:

Please don't go down the Yarm (Old Xarm) for right now, or if you do, please be very careful.  Kiwamu and I are set up to take beam scan measurements down the walkway, and so there are some cables / carts / other stuff down there.  We are going to get dinner really quickly before beginning the measurements. 

Right now, the PSL shutter is Closed, so there is no beam hazard outside of the chambers, just crowded space hazard.

 

  2961   Thu May 20 20:03:37 2010 JenneUpdateSAFETYDon't walk down the Y(new convention) arm tonight!

Please don't go down the Yarm (Old Xarm) for right now, or if you do, please be very careful.  Kiwamu and I are set up to take beam scan measurements down the walkway, and so there are some cables / carts / other stuff down there.  We are going to get dinner really quickly before beginning the measurements. 

Right now, the PSL shutter is Closed, so there is no beam hazard outside of the chambers, just crowded space hazard.

  2960   Thu May 20 14:18:59 2010 kiwamuUpdateGreen Lockingmode profile of 40m cavity

The mode profile of the green beam going through 40m cavity was measured.

According to the fitting the coupling efficiency to the cavity is 98.46%, but still the beam looks loosely focused.

This measurement has been done by using the oplev legs (entry #2957) to allow the beam to go through the 40m walkway.

With a beam scan set on a movable cabinet, I measured it along the 40m chamber.

Since the plot looks not so nice,  I am going to work on this measurement a little bit more after I improve the mode matching.

 


Here is the parameters from the fitting

target waist [mm] 2.662
measured waist x [mm] 2.839
measured waist x [mm] 3.111
   
target waist position [m] 43.498
measured waist position x [m] 42.579
measured waist position y [m] 38.351

 

I believe the error for the travel length was within 0.5 meter. The length was always measured by a tape measure.

A thing I found was that: spatial jittering of the beam gets bigger as the beam goes further. This is the main source of the error bar for the spot size. 

 

 

Attachment 1: MMT40mcavity.png
MMT40mcavity.png
  2959   Thu May 20 13:29:40 2010 kiwamuUpdateGreen Lockingmode profile at PPKTP crystal

 I measured again the mode profile of the beam going through the PPKTP crystal by using the beam scan.

The aimed beam waist is 50 um (as described in entry 2735),

and the measured profile had pretty good waist of wx=51.36 +/- 0.0999 um and wy=49.5 +/- 0.146 um 

The next things I have to do are - (1). re-optimization of the temperature of the crystal (2). measurement of the conversion efficiency

The attached figure is the result of the measurement. 

Attachment 1: PPKTPmode.png
PPKTPmode.png
  2958   Thu May 20 13:12:28 2010 josephbUpdateCDSPreparations for testing lsc,lsp, scy,spy together

In /cvs/cds/caltech/target/fb modified:

master: cleaned up so only io1 (IO processor), LSC, LSP, SCY, SPY were listed, along with their associated tpchan files.

daqdrc: fixed "dcu_rate 9 = 32768" to "dcu_rate 9 = 65536" (since the IO processor is running at 64k)

Added "dcu_rate 21 = 16384" and "dcu_rate 22 = 16384"

Changed "set gds_server = "megatron" "megatron" "megatron" 9 "megatron" 10 "megatron" 11;" to

set gds_server = "megatron" "megatron" 9 9;

The above change was made after reading Rolf's Admin guide: http://lhocds.ligo-wa.caltech.edu:8000/40m/Upgrade_09/CDS?action=AttachFile&do=get&target=RCG_admin_guide.pdf
The set gds_server is simply telling which computer the gds daemons are running on, and we don't need to do it 5 times.


In /cvs/cds/caltech/gds/params modified:

testpoint.par: added C-node7 and C-node8 for SCY and SPY respectively.

  2957   Thu May 20 12:34:46 2010 kiwamuConfiguration40m Upgradingoptical breadboards with legs

Yesterday Steve and I revived two legs to mount some optical breadboards outside of the end table.

These legs had been used as oplev's mounts many years ago, but now they are served for 40m upgrading. These are really nice.

By putting them on the side of the end table, a mirror mounted on the top of the leg can reflect the beam outside of the end table.

Once we pick off the green beam from the end table to its outside, the green beam can propagate through the 40m walkway along the Y-arm.

So that we can measure the beam profile as it propagates.

These legs are also going to be used during mode matching of the vacuum optics.

Attachment 1: leg1_small.png
leg1_small.png
Attachment 2: leg2_small.png
leg2_small.png
  2956   Thu May 20 12:10:44 2010 kiwamuUpdatePhotosETMY end table

 I updated the photo of ETMY end table on the wiki.

http://lhocds.ligo-wa.caltech.edu:8000/40m/Optical_Tables

Attachment 1: ETMY_s.png
ETMY_s.png
  2955   Thu May 20 10:06:56 2010 AidanHowToPhase CameraPhase Camera- raw data video


 

  2954   Wed May 19 22:28:05 2010 KojiUpdateIOOHooray! We locked the MC! (and some other stuff)

Good! What was the key?

The MC2 spot looks very high, but don't believe the TV image. Believe the result of script/A2L/A2L_MC2. What you are looking at is the comparison of the spot at the front surface and the OSEMs behind the mirror.

Quote:

[Jenne, Kevin]

We opened up the MC chambers again, and successfully got the MC locked today!  Hooray!  This meant that we could start doing other stuff....

First, we clamped the Faraday.  I used the dog clamps that Zach left wrapped in foil on the clean cart.  I checked with a card, and we were still getting the 00 mode through, and I couldn't see any clipping.  2 thumbs up to that.

Then we removed the weight that was on the OMC table, in the way of where MMT2 needs to go.  We checked the alignment of the MC, and it still locks on TEM00, but the spot looks pretty high on MC2 (looking at the TV view). We're going to have to relevel the table when we've got the MMT2 optic in the correct place.

We were going to start moving the PZT steering mirror from the BS table to the IOO table, place MMT2 on the OMC table, and put in a flat mirror on the BS table to get the beam out to the BS oplev table, but Steve kicked us out of the chambers because the particle count got crazy high.  It was ~25,000 which is way too high to be working in the chambers (according to Steve).  So we closed up for the day, and we'll carry on tomorrow.  

Photos of the weight before we removed it from the OMC table, and a few pictures of the PZT connectors are on Picasa

 

  2953   Wed May 19 16:09:11 2010 josephbUpdateCDSRacks to small for IO Chassis rails

So I discovered the hard way that the racks are not standard width, when I was unable to place a new IO chassis into the racks with rails attached.  The IO chassis is narrow enough to fit through without the rails however. 

I've talked to Steve and we decided on having some shelves made.  I've asked Steve to get us 6.  1 for each end (2), 1 for SUS, 1 for LSC, 1 for IO, and 1 extra.

  2952   Wed May 19 16:00:18 2010 JenneUpdateIOOHooray! We locked the MC! (and some other stuff)

[Jenne, Kevin]

We opened up the MC chambers again, and successfully got the MC locked today!  Hooray!  This meant that we could start doing other stuff....

First, we clamped the Faraday.  I used the dog clamps that Zach left wrapped in foil on the clean cart.  I checked with a card, and we were still getting the 00 mode through, and I couldn't see any clipping.  2 thumbs up to that.

Then we removed the weight that was on the OMC table, in the way of where MMT2 needs to go.  We checked the alignment of the MC, and it still locks on TEM00, but the spot looks pretty high on MC2 (looking at the TV view). We're going to have to relevel the table when we've got the MMT2 optic in the correct place.

We were going to start moving the PZT steering mirror from the BS table to the IOO table, place MMT2 on the OMC table, and put in a flat mirror on the BS table to get the beam out to the BS oplev table, but Steve kicked us out of the chambers because the particle count got crazy high.  It was ~25,000 which is way too high to be working in the chambers (according to Steve).  So we closed up for the day, and we'll carry on tomorrow. 

 

Photos of the weight before we removed it from the OMC table, and a few pictures of the PZT connectors are on Picasa

  2951   Wed May 19 14:36:46 2010 AidanHowToPhase CameraPhase Camera algorithm and stuff

 I had a think about the algorithm we might use for the phase camera measurement. MATLAB has an fft function that will allow us to extract the data that we need with a single command.

We record a series of images from a camera and put them into a 3D array or movie, image_arr, where the array parameters are [x-position, y-position, time], i.e. a 2D slice is a single frame from the camera. Then we can do an FFT on that object with the syntax, f3D = fft(image_arr, [ ], 3), which only does the FFT on the temporal components. The resulting object is a 3D array where each 2D slice is an 2D array of amplitude and phase information across the image for a single temporal frequency of the movie.

So if we recorded a movie for 1s where the sample rate is 58Hz, then the 1st frame of f3D is just a DC image of the movie, the 2nd frame are the complex 1Hz components of the movie, etc all the way up to 29Hz. 

Suppose then that we have a image, part of which is being modulated, e.g. a chopper wheel rotating at 20 or 24Hz, or a laser beam profile which contains a 1kHz beat between a sideband and a reference beam. All we have to do is sample at at least twice that modulation frequency, run the command in MATLAB, and then we immediately get an image which contains the phase and magnitude information that we're interested in (in the appropriate 2D slice o the FFT).

As an example, I recorded 58 frames of data from a camera, sampling at 58Hz, which was looking at a spinning chopper wheel. There was a white sheet of paper behind the wheel which was illuminated from behind by a flashlight. The outer ring was chopping at 24Hz and the inner ring was chopping at 20Hz. I stuck all the images into the 3D array in MATLAB, did the transformation and picked out the DC, 20Hz and 24Hz signals. The results are shown in the attached PDFs which are:

  1. phase_camera_DC_comp.pdf - a single image from the camera and the DC component (zoomed in) of the FFT
  2. phase_camera_F1_comp.pdf - the magnitude and phase information of the 20Hz component of the FFT
  3. phase_camera_F2_comp.pdf - the magnitude and phase information of the 24Hz component of the FFT (this PDF contains a typo that says 25Hz).
  4. load_raw_data.m - the MATLAB routine that loads the saved data from the camera and does the FFT

You can, and I have, run the MATLAB engine from C directly. This will allow you to transfer the data from the camera to MATLAB directly in memory, rather than via the disk, but it does need proper memory allocation to avoid segmentation faults - that was too frustrating for me in the short term. In this case, the 58 frames were recorded to a file as a contiguous block of data which I then loaded into MATLAB, so it was slower than it might've otherwise been. Also the computer I was running this on was a bit of a clunker so it took a bit of time to do the FFT.

The data rate from the camera was 58fps x (1024 x 1024) pixels per frame x 2 bytes per pixel = 116MB per second. If we were to use this technique in a LIGO phase camera, where we want to measure a modulation which is around 1kHz, then we'd need a sample rate of at least 2kHz, so we're looking at at least a 30x reduction in the resolution. This is okay though - the original phase camera had only ~4000 spatial samples. So we could use, for instance, the Dalsa Falcon VGA300 HG which can give 2000 frames per second when the region of interest is limited to 64 pixels high.

Attachment 1: phase_camera_DC_comp.pdf
phase_camera_DC_comp.pdf
Attachment 2: phase_camera_F1_comp.pdf
phase_camera_F1_comp.pdf
Attachment 3: phase_camera_F2_comp.pdf
phase_camera_F2_comp.pdf
Attachment 4: load_raw_data.m
% load a raw data file into MATLAB

fid = fopen('phase_camera_data.dat');
n1 = 750;
A3D = ones(n1, n1, 58);

for jj = 1:58
    A = fread(fid, [1024, 1024], 'uint16');
    A3D(:,:,jj) = A((512-floor(n1/2)):(512-floor(n1/2))+n1-1, ...
                    (512-floor(n1/2)):(512-floor(n1/2))+n1-1);
... 64 more lines ...
  2950   Tue May 18 23:03:08 2010 JenneUpdateIOONo real progress....

[Jenne, Kevin]

No real progress today.  We opened the chambers and again tried to lock the MC.  Gave up after ~2.5 hours (and closed up the chambers with light doors, replaced manual beam block, etc...).  With Koji's helpful coaching, hopefully we'll finally get it done tomorrow.  Then we can move forward with the actual to-do list. 

 

  2949   Tue May 18 16:44:35 2010 KojiUpdateIOOFirst steps toward MC mode measuring

Here is the upadted list http://lhocds.ligo-wa.caltech.edu:8000/40m/Upgrade_09/Optics

Quote:

I will update how the mirrors should be migrated from the table to the table. 

 

  2948   Tue May 18 16:19:19 2010 josephbUpdateCDSWe have two new IO chassis

We have 2 new IO chassis with mounting rails and necessary boards for communicating to the computers.  Still need boards to talk to the ADCs, DACs, etc, but its a start.  These two IO chassis are currently in the lab, but not in their racks.

They will installed into 1X4 and 1Y5 tomorrow.  In addition to the boards, we need some cables, and the computers need the approriate real time operating systems setup.  I'm hoping to get Alex over sometime this week to help work on that.

ELOG V3.1.3-